OpenStudy (elisaneedshelp):

How do i simplify sqrt85? I only reached sqrt17 times sqrt5 and i got stuck Please help

OpenStudy (elisaneedshelp):

@nincompoop

OpenStudy (elisaneedshelp):

@rvc

OpenStudy (elisaneedshelp):

@lewopz

rvc (rvc):

5sqrt17

OpenStudy (whpalmer4):

\[85=5*17\]Both those numbers are prime, so you cannot simplify \(\sqrt{85}\).

OpenStudy (anonymous):

what he said^^

OpenStudy (whpalmer4):

@rvc \[5\sqrt{17} = \sqrt{25*17}\]

OpenStudy (elisaneedshelp):

So would the answer just be to leave it like sqrt85 or 5sqrt17?

rvc (rvc):

yes

OpenStudy (anonymous):

5sqrt17 is wrong

OpenStudy (anonymous):

so with leave as is or plug into a calculator

OpenStudy (elisaneedshelp):

my calculator agrees with you @lewopz

rvc (rvc):

without using calculator

rvc (rvc):

is this correct @lewopz

OpenStudy (anonymous):

what is?

rvc (rvc):

the ans

OpenStudy (elisaneedshelp):

For sqrt576 i started off with sqrt4 times sqrt144 what do i do next?

rvc (rvc):

24 is the ans

OpenStudy (anonymous):